# Neural Graduated Assignment for Maximum Common Edge Subgraphs

### v1

@@ Abstract
The Maximum Common Edge Subgraph (MCES) problem is a crucial challenge with significant implications in domains such as biology and chemistry. Traditional approaches, which include transformations into max-clique and search-based algorithms, suffer from scalability issues when dealing with larger instances. This paper introduces “Neural Graduated Assignment” (NGA), a simple, scalable, unsupervised-training-based method that addresses these limitations by drawing inspiration from the classical Graduated Assignment (GA) technique. Central to NGA is stacking of neural components that closely resemble the GA process, but with the reparameterization of learnable temperature into higher dimension. We further theoretically analyze the learning dynamics of NGA, showing its design leads to fast convergence, better exploration-exploitation tradeoff, and ability to escape local optima. Extensive experiments across MCES computation, graph similarity estimation, and graph retrieval tasks reveal that NGA not only significantly improves computation time and scalability on large instances but also enhances performance compared to existing methodologies. The introduction of NGA marks a significant advancement in the computation of MCES and offers insights into other assignment problems.
@@ 1 Introduction
The Maximum Common Edge Subgraph (MCES) problem is a cornerstone task in combinatorial optimization [30], particularly significant within the realms of biology and chemistry [11]. As a variant of the broader Maximal Common Subgraph (MCS) problem [6], MCES stands alongside the Maximal Common Induced Subgraph (MCIS) in its complexity and utility. Efficiently solving the MCES problem at scale is of both theoretical and practical significance in this context. For instance, in drug discovery, identifying maximal common substructures between molecules can reveal shared pharmacological properties. In cybersecurity, comparing network traffic graphs via MCES enables the detection of recurring attack patterns [11]. Established approaches for solving MCES have been successfully integrated into widely-used cheminformatics libraries such as RDKit [4] and Molassembler [39], which have become indispensable in both industrial and academic research.
@@
Figure 1: For two labeled graphs, e.g. molecules, the MCES is highlighted in circle and the node correspondences are encoded in arrowed lines.
@@
The MCES problem involves identifying a subgraph that contains the maximum number of edges common to both input graphs. An example of MCES is illustrated in Figure 1. It is inherently NP-complete [14], posing significant computational challenges in terms of scalability and efficiency. Historically, it has been tackled through transformations into the maximum clique problem [5] or search-based branch-and-bound algorithms [34, 27, 26]. However, these traditional methods often struggle with large-scale graphs: transformations can introduce additional computational overhead, while search-based approaches suffer from exponential scaling, rendering them impractical for complex instances. This challenging landscape underscores the need for novel techniques that can deliver efficient and reliable solutions to the MCES problem without compromising accuracy.
@@
To address the challenges of MCES, we propose Neural Graduated Assignment (NGA) – a novel neuralized optimization framework designed to approximate MCES solutions in polynomial time. Inspired by principles from statistical physics, particularly the classic Graduated Assignment (GA) method [32], our approach seeks to find an assignment matrix that identifies node correspondences between the two graphs, such that the number of preserved edges – i.e., edges that exist between matched node pairs in both graphs – is maximized. NGA tackles the aforementioned intrinsic challenges in the MCES problem by: (1) constructing an Association Common Graph (ACG) and formulating MCES as a Quadratic Assignment Problem (QAP), which ensures the extraction of exact common subgraphs between input graphs; (2) iteratively updating the learned assignments with the reparameterization of the temperature in GA into higher dimension; and (3) optimizing this formulation via unsupervised training [^1] Unsupervised training means that no information about the MCES itself—such as what the MCES is or the MCES size—is required during training, including in the loss function [37]., thereby eliminating the need for training data. During optimization, the learned assignments correspond to a current identified common edge subgraph that serves as a lower bound, which is progressively improved toward the optimal MCES. An illustrative example of such improved lower bound is provided in Figure 12 of the Appendix.
@@
Despite the simplicity, we theoretically justify the superiority of NGA by showing: (1) how NGA updates around local optima; (2) the implicit exploration-exploitation mechanism impacts the optimization trajectory; and (3) the accelerating effect on the convergence. By embedding these operations within a neural context, NGA achieves what GA could not: scalability, adaptability, and seamless integration with machine learning frameworks. This theoretical underpinning lays the groundwork for a new class of algorithms capable of rethinking assignment problems from a neural perspective.
@@
To validate the efficacy of our approach, we conduct a series of challenging experiments across various settings, including large-scale MCES computation, graph similarity estimation, and structure-based graph retrieval. The results consistently demonstrate the superior efficiency and performance of our method compared to existing ones. These findings not only highlight the potential of our approach in advancing MCES computation but also suggest its applicability to diverse domains. We summarize our contributions as follows:
@@
  • A novel non-search-based approach. We propose NGA – the first neural-style algorithm to approximate the MCES solution efficiently in polynomial time without relying on exhaustive exploration of the solution space.
@@
  • Unsupervised framework. NGA operates in a fully unsupervised manner, eliminating the need for supervision signals. In cases where enumeration or search of the exact solution becomes computationally infeasible, our approach provides efficient approximations.
  • Theoretical analysis. We present a theoretical analysis of NGA about its behavior centered at local optima, and its implicit capacity for both exploration and exploitation.
@@ 3 Preliminary
Two graphs, $G_{1}$ and $G_{2}$, are isomorphic if there exists a bijective (one-to-one) mapping between their nodes such that any two nodes in $G_{1}$ are connected by an edge if and only if their corresponding images in $G_{2}$ are also connected. A common subgraph of two graphs $G_{1}$ and $G_{2}$ is a graph $G_{12}$ that is isomorphic to a subgraph of $G_{1}$ and a subgraph of $G_{2}$. Although there are possibly many common subgraphs between two graphs, our focus will be on the MCES [1], which is a subgraph with the maximal number of edges common to both $G_{1}$ and $G_{2}$.
Graduated Assignment.
GA is a classical optimization method rooted in principles of statistical physics and commonly applied to assignment problems such as graph matching [16, 8]. GA operates by iteratively updating a soft assignment matrix, which represents the probabilistic relationship between elements of two sets (e.g., nodes, edges) that are being matched. Given two graphs $G_{1}$ and $G_{2}$, the goal is to find a hard assignment matrix $\mathbf{P}\in\{0,1\}^{n_{1}\times n_{2}}$ that maximizes a compatibility function while adhering to row and column constraints. The optimization problem is formalized as:
@@
where $\mathbf{A}$ is the affinity matrix derived from the structural information of $G_{1}$ and $G_{2}$ and $\mathbf{1}_{n}$ is a column vector of length $n$ whose elements are all equal to 1. GA relaxes $\mathbf{P}$ into a soft assignment matrix $\mathbf{S}\in[0,1]^{n_{1}\times n_{2}}$, allowing continuous values, and optimizes the following relaxed objective:
$\max_{\mathbf{S}}\;\;\mathrm{vec}(\mathbf{S})^{\top}\mathbf{A}\mathrm{vec}(\mathbf{S})-\frac{1}{\beta}\mathbb{H}(\mathbf{S}),$ (2)
where $\mathbb{H}(\mathbf{S})$ measures the entropy of $\mathbf{S}$. The second entropic regularization term encourages the solution to remain smooth (i.e., probabilistic). Rooted in the KKT conditions, GA can be briefly summarized as in Alg. 3 in Appendix A. Readers are referred to [33, 32] for more details of GA.
@@ 4.2.1 ACG for Learning the Correspondences
Since only the common edges are included in the ACG, Proposition 1 shows that we can always extract common subgraphs from the constructed association common subgraph. Detailed illustration of Proposition 1 can be found in Appendix B. Considering the adjacency matrix $\mathcal{A}_{\Diamond}$ of $G_{1}\Diamond G_{2}$ as the affinity matrix, we can then give the QAP formulation of objective $J$ for MCES as:
@@ 4.2.2 Neural Graduated Assignment
Since many nodes have the same labels in molecular graphs, the initial assignment is a soft probability distribution where many entries have moderate probabilities, which leaves significant ambiguity in the assignments. GA is a feasible approach to improve the initial solution because it can gradually reduce the energy of the objective function. GA relies on a pre-defined temperature parameter $\beta$ in its optimization formulations to iteratively refine solutions by controlling the scale of probabilities or soft assignment [33, 16]. While GA treat $\beta$ as a static or manually scheduled hyperparameter, it suffers from two key limitations: (1) Manual Tuning Overhead: Selecting an optimal schedule or fixed value for $\beta$ often requires exhaustive tuning, depending on the specific task or dataset. (2) Limited Adaptability: A fixed or scheduled $\beta$ is static and cannot dynamically adapt to the complexity of individual problems or local structures within data.
@@ 4.2.3 Gumbel Sampling for Optimization
[…]
@@
[…]
@@
where $g$ is sampled from a standard Gumbel distribution with the cumulative distribution function. The Gumbel term models the distribution of extreme values derived from another distribution. With Eq. (8), we are able to sample repeatedly a batch of assignment matrices from the original distribution in a differentiable way. This property actively benefits solution space exploration and makes it easier to find the optimal solution. These sampled assignment matrices are discretized by Hungarian algorithm and evaluated by Eq. (7). The best-performing solution among them is chosen as the final solution. The balance between exploration and speed can be adjusted by tuning the number of Gumbel samples.
@@ 5 Theoretical Analysis
In Alg. 1, let $\beta_{l}=\mathbf{W}_{1}^{(l)\top}\mathbf{W}_{2}^{(l)}$, then we define two phases according to the value of $\beta_{l}$: 1. Exploration phase: =$\mathcal{E}=\{l|\beta_{l}<0\}$; 2. Exploitation phase: $\mathcal{S}=\{l|\beta_{l}>0\}$.
Appendix D.1 provides a proof explaining why the sign of $\beta_{l}$ represents these two processes, respectively. Imposing a negative and learnable $\beta_{l}$ in NGA grants the model the ability to balance exploration and exploitation. This contrasts with the traditional GA algorithm, which relies on a monotonically increasing positive $\beta$ parameter. As $\beta$ always increases in GA, the algorithm lacks exploration capabilities and is more prone to getting stuck in local optima. By allowing negative values, NGA effectively incorporates exploration into the optimization, avoiding premature convergence and enabling the identification of better solutions.
@@ 6.1 Experimental Setup
The MCES problem is NP-complete [14, 34], meaning no known polynomial-time solutions exist. While search-based algorithms can find exact solutions with enough computational resources, they quickly become infeasible for large graphs. Non-search-based methods aim to approximate solutions efficiently without exhaustive search. We compare our method with search-based RASCAL [34] and Mcsplit [27], and use GA as a baseline since we formulate MCES as a QAP. We also adapt the supervised learning method NGM [40] for comparison. For the graph similarity and retrieval tasks, we compare with state-of-the-art baselines: NeuroMatch [25], SimGNN [2], GMN [24], XMCS [35], and INFMCS [23]. Appendix C.3 details baseline choices and configurations.
For fair comparisons, supervised learning approaches are trained for a total of 200 epochs. This duration is selected to provide sufficient learning time for the models to converge, ensuring robust evaluation of their performance. For unsupervised learning and search methods, we follow previous works [3] to set a practical time budget of 60 seconds for each pair of graphs. This constraint reflects real-world scenarios where computation time is limited and must be optimized for efficiency. Note that while search methods are capable of identifying exact solutions given unlimited computational resources, this assumption is impractical for handling large graph pairs in real-world applications.
@@
As for the graph similarity computation, the Johnson similarity [20] is calculated according to
@@ 6.2 Solving MCES Problems
The key property of NGA is its ability to approximate solutions in polynominal time complexity. As shown in Table 6.2, our model outperforms baselines on all datasets. Our results are very close to optimal in all datasets, which shows that NGA can effectively approximate solutions within limited time. The supervised graph matching method does not perform well, possibly because the MCES problem is inherently multi-modal, meaning it has multiple equally optimal solutions. It is thus difficult for supervised learning method to learn some patterns from multiple ground truth node matching signals. The search method, on the other hand, suffers from scalability issues due to the NP-complete nature of the MCES problem. In the following part, we provide several case studies that investigate the effects of increasing the time budget on performance.
We selected several cases for detailed analysis. The largest MCES solution found thus far as time grows is illustrated in Figure LABEL:fig:best_solution_so_far (in Appendix G.1). The visualizations of the MCES structures identified by our NGA method and the second-best method, RASCAL, are shown in Figure 3 and Figure 13 (in Appendix G.1). From these figures, it is evident that our method can find better solutions in less time, demonstrating its superiority over other competitive baselines.
@@
Table 2: Results of MSE (×$\times 10^{-3}$) $\downarrow$ for graph similarity.

### v2

@@ Abstract
The Maximum Common Edge Subgraph (MCES) problem is a crucial challenge with significant implications in domains such as biology and chemistry. Traditional approaches, which include transformations into max-clique and search-based algorithms, suffer from scalability issues when dealing with larger instances. This paper introduces “Neural Graduated Assignment” (NGA), a simple, scalable, unsupervised-training-based method that addresses these limitations. Central to NGA is stacking of differentiable assignment optimization with neural components, enabling high-dimensional parameterization of the matching process through a learnable temperature mechanism. We further theoretically analyze the learning dynamics of NGA, showing its design leads to fast convergence, better exploration-exploitation tradeoff, and ability to escape local optima. Extensive experiments across MCES computation, graph similarity estimation, and graph retrieval tasks reveal that NGA not only significantly improves computation time and scalability on large instances but also enhances performance compared to existing methodologies. The introduction of NGA marks a significant advancement in the computation of MCES and offers insights into other assignment problems.
@@ 1 Introduction
Background. The Maximum Common Edge Subgraph (MCES) problem is a cornerstone task in combinatorial optimization [34], particularly significant within the realms of biology and chemistry [12]. As a variant of the broader Maximal Common Subgraph (MCS) problem [6], MCES stands alongside the Maximal Common Induced Subgraph (MCIS) in its complexity and utility. Efficiently solving the MCES problem at scale is of both theoretical and practical significance in this context. For instance, in drug discovery, identifying maximal common substructures between molecules can reveal shared pharmacological properties. In cybersecurity, comparing network traffic graphs via MCES enables the detection of recurring attack patterns [12]. Established approaches for solving MCES have been successfully integrated into widely-used cheminformatics libraries such as RDKit [4] and Molassembler [43], which have become indispensable in both industrial and academic research.
@@
Figure 1: For two labeled graphs (molecules), the maximum common edge subgraph (MCES) is highlighted in circle and the node correspondences are encoded in arrowed lines.
@@
Problem & Existing Challenges. MCES involves identifying a subgraph that contains the maximum number of edges common to both input graphs. An example is in Figure 1. It is inherently NP-complete [16], posing significant computational challenges in terms of scalability and efficiency. Historically, it has been tackled through transformations into the maximum clique problem [5] or search-based branch-and-bound algorithms [38, 31, 30]. However, these traditional methods often struggle with large-scale graphs: transformations can introduce additional computational overhead, while search-based approaches suffer from exponential scaling, rendering them impractical for complex instances. This challenging landscape underscores the need for novel techniques that can deliver efficient and reliable solutions to the MCES problem without compromising accuracy.
@@
Our approach. To address the challenges of MCES, we propose Neural Graduated Assignment (NGA) – a novel neuralized optimization framework designed to approximate MCES solutions in polynomial time. Inspired by annealing mechanism in statistical physics [18], our approach seeks to find an assignment matrix that identifies node correspondences between the two graphs, such that the number of preserved edges – i.e., edges that exist between matched node pairs in both graphs – is maximized. NGA tackles the aforementioned intrinsic challenges in the MCES problem by: (1) constructing an Association Common Graph (ACG) and formulating MCES as a Quadratic Assignment Problem (QAP), which ensures the extraction of exact common subgraphs between input graphs; (2) iteratively updating the learned assignments through a high-dimensional and learnable temperature parameterization; and (3) optimizing this formulation via unsupervised training [^1] Unsupervised training means that no information about the MCES itself—such as what the MCES is or the MCES size—is required during training, including in the loss function [41]., thereby eliminating the need for training data. During optimization, the learned assignments correspond to a current identified common edge subgraph that serves as a lower bound, which is progressively and monotontically improved toward the optimal MCES. An illustrative example of such improved lower bound is provided in Figure 12 of the Appendix.
@@
Theory. Despite the simplicity, we theoretically justify the superiority of NGA by showing: (1) how NGA updates around local optima; (2) the implicit exploration-exploitation mechanism impacts the optimization trajectory; and (3) the accelerating effect on the convergence. By embedding these operations within a neural context, NGA achieves strong scalability, adaptability, and seamless integration with machine learning frameworks. This theoretical underpinning lays the groundwork for a new class of algorithms capable of rethinking assignment problems from a neural perspective.
@@
Results & Contribution. To validate the efficacy of our approach, we conduct a series of challenging experiments across various settings, including large-scale MCES computation, graph similarity estimation, and structure-based graph retrieval. The results consistently demonstrate the superior efficiency and performance of our method compared to existing ones. These findings not only highlight the potential of our approach in advancing MCES computation but also suggest its applicability to diverse domains. We summarize our contributions as follows:
@@
  • A novel and strong approach. We for the first time formulate the MCES problem via the construction of an ACG, and based on this formulation, we propose NGA – the first neural-style algorithm to approximate the MCES solution efficiently in polynomial time without relying on exhaustive exploration of the solution space, delivering superior performance.
@@
  • Training-data-free. NGA operates in a fully unsupervised manner, eliminating the need for supervision signals. In cases where enumeration or search of the exact solution becomes computationally infeasible, our approach provides efficient approximations.
  • Theoretical analysis. We provide the first theoretical analysis on the behavior of dynamic and parameterized temperature in NGA, shedding light on its behavior at local optima, convergence properties and adaptive dynamics.
@@ 3 Preliminary
Two graphs, $G_{1}$ and $G_{2}$, are isomorphic if there exists a bijective mapping between their nodes such that any two nodes in $G_{1}$ are connected by an edge if and only if their corresponding images in $G_{2}$ are also connected. A common subgraph of two graphs $G_{1}$ and $G_{2}$ is a graph $G_{12}$ that is isomorphic to a subgraph of $G_{1}$ and a subgraph of $G_{2}$. Although there are possibly many common subgraphs between two graphs, our focus will be on the MCES [1], which is a subgraph with the maximal number of edges common to both $G_{1}$ and $G_{2}$.
Quadratic Assignment Problem.
Given two graphs $G_{1}$ and $G_{2}$, the goal is to find a hard assignment matrix $\mathbf{P}\in\{0,1\}^{n_{1}\times n_{2}}$ that maximizes a compatibility function while adhering to row and column constraints. The optimization problem is formalized as:
@@
where $\mathbf{A}$ is the affinity matrix derived from the structural information of $G_{1}$ and $G_{2}$ and $\mathbf{1}_{n}$ is a column vector of length $n$ whose elements are all equal to 1. A common practice is to relax $\mathbf{P}$ into a soft assignment matrix $\mathbf{S}\in[0,1]^{n_{1}\times n_{2}}$, allowing continuous values, and optimizes the following relaxed objective:
$\max_{\mathbf{S}}\;\;\mathrm{vec}(\mathbf{S})^{\top}\mathbf{A}\mathrm{vec}(\mathbf{S})$ (2)
@@ 4.2.1 ACG for Learning the Correspondences
Since only the common edges are included in the ACG, Proposition 1 shows that we can always extract common subgraphs from the constructed association common subgraph. Formal proof of Proposition 1 can be found in Appendix B. Considering the adjacency matrix $\mathcal{A}_{\Diamond}$ of $G_{1}\Diamond G_{2}$ as the affinity matrix, we can then give the QAP formulation of objective $J$ for MCES as:
@@ 4.2.2 Neural Graduated Assignment
Since many nodes have the same labels in molecular graphs, the initial assignment is a soft probability distribution where many entries have moderate probabilities, which leaves significant ambiguity in the assignments. Traditional assignment-based optimization methods often rely on a predefined temperature parameter $\beta$ to iteratively refine solutions by controlling the scale of probabilities or soft assignments. However, using a fixed or manually scheduled temperature introduces two key limitations: (1) Manual Tuning Overhead: Selecting an optimal schedule or fixed value for $\beta$ often requires exhaustive tuning, depending on the specific task or dataset. (2) Limited Adaptability: A fixed or scheduled $\beta$ is static and cannot dynamically adapt to the complexity of individual problems or local structures within data.
@@ 4.2.3 Gumbel Sampling for Optimization
Since $\mathbf{S}$ is a relaxed from $\mathbf{P}$, the Hungarian algorithm [7] is commonly employed as a deterministic post-processing step to bridge the gap between them, i.e. $\mathbf{P}=\mathrm{Hungarian}(\mathbf{S})$. The adjacency matrix of predicted common subgraph $\mathcal{A}{\textsubscript{pred}}$ is obtained by:
$\mathcal{A}{\textsubscript{pred}}=\left(\mathrm{vec}(\mathbf{P})\mathrm{vec}(\mathbf{P})^{\top}\right)\odot\mathcal{A}_{\Diamond},$ (7)
@@
From a probabilistic perspective, $\mathbf{S}$ represents a latent distribution of assignment matrices. Assignment with the highest probability is selected by Hungarian algorithm. However, there may be better solutions within the distribution, especially when the quality of solutions can be easily assessed. Therefore, we switch to Gumbel-Sinkhorn [32] by substituting Eq. (6) with
@@
where $g$ is sampled from a standard Gumbel distribution with the cumulative distribution function. The Gumbel term models the distribution of extreme values derived from another distribution. With Eq. (8), we can sample repeatedly a batch of assignment matrices from the original distribution in a differentiable way. This property actively benefits solution space exploration and makes it easier to find the optimal solution. These sampled assignment matrices are discretized by Hungarian algorithm and evaluated by Eq. (7). The best-performing solution among them is chosen as the final solution. The balance between exploration and speed can be adjusted by tuning the number of Gumbel samples.
@@ 5 Theoretical Analysis
In Alg. 1, let $\beta_{l}=\mathbf{W}_{1}^{(l)\top}\mathbf{W}_{2}^{(l)}$, then we define two phases according to the value of $\beta_{l}$: 1. Exploration phase: $\mathcal{E}=\{l|\beta_{l}<0\}$; 2. Exploitation phase: $\mathcal{S}=\{l|\beta_{l}>0\}$.
Appendix D.1 provides a proof explaining why the sign of $\beta_{l}$ represents these two processes, respectively. Imposing a negative and learnable $\beta_{l}$ in NGA grants the model the ability to balance exploration and exploitation. By allowing the two phases, NGA effectively incorporates exploration and exploitation into the optimization, avoiding premature convergence and enabling the identification of better solutions.
@@ 6.1 Experimental Setup
The MCES problem is NP-complete [16, 38], meaning no known polynomial-time solutions exist. While search-based algorithms can find exact solutions with enough computational resources, they quickly become infeasible for large graphs. Non-search-based methods aim to approximate solutions efficiently without exhaustive search. We compare our method with search-based RASCAL [38] and Mcsplit [31], and use Graduated Assignment (GA) [37, 18] and Gurobi solver as baselines since we can formulate MCES as a QAP. Similar to these solvers, our NGA tries to solve each MCES instance case by case. We also adapt the supervised learning method NGM [45] for comparison. For the graph similarity and retrieval tasks, we compare with state-of-the-art baselines: NeuroMatch [28], SimGNN [2], GMN [27], XMCS [39], and INFMCS [25]. Appendix C.3 details baseline choices and configurations.
For fair comparisons, supervised learning approaches are trained for a total of 200 epochs. This duration is selected to provide sufficient learning time for the models to converge, ensuring robust evaluation of their performance. For unsupervised learning and search methods, we follow previous works [3] to set a practical time budget of 60 seconds for each instance. This constraint reflects real-world scenarios where computation time is limited and must be optimized for efficiency. While search methods are capable of identifying exact solutions given unlimited computational resources, this assumption is impractical for handling large graph pairs in real-world applications.
@@
As for the graph similarity computation, the Johnson similarity [22] is calculated as
@@ 6.2 Solving MCES Problems
The key property of NGA is its ability to approximate solutions in polynominal time complexity. As shown in Table 6.2, our model outperforms baselines on all datasets. Our results are very close to optimal in all datasets, which shows that NGA can effectively approximate solutions within limited time. We have additionally analyzed how often the MCES found by NGA matches the optimal solution in Appendix G.1. The supervised graph matching method does not perform well, possibly because the MCES problem is inherently multi-modal, meaning it has multiple equally optimal solutions. It is thus difficult for supervised learning method to learn some patterns from multiple ground truth node matching signals. The search method, on the other hand, suffers from scalability issues due to the NP-complete nature of the MCES problem. In the following part, we provide several case studies that investigate the effects of increasing the time budget on performance.
We selected several cases for detailed analysis. The largest MCES solution found thus far as time grows is illustrated in Figure 12(a) (in Appendix G.2). The visualizations of the MCES structures identified by our NGA method and the second-best method, RASCAL, are shown in Figure 3 and Figure 13 (in Appendix G.2). From these figures, it is evident that our method can find better solutions in less time, demonstrating its superiority over other competitive baselines.
@@
Table 2: Results of MSE ($\times 10^{-3}$) $\downarrow$ for graph similarity.